OpenStudy (anonymous):

Find the surface area of a conical grain storage tank that has a height of 42 meters and a diameter of 20 meters. Round the answer to the nearest square meter. 3,028 m2 3,971 m2 1,357 m2 1,671 m2

OpenStudy (anonymous):

use this formula http://math.about.com/od/formulas/ss/surfaceareavol_2.htm radious is half of diamter

OpenStudy (anonymous):

pi*r*s to find s use pythogorean theorom

OpenStudy (anonymous):

sqrt(10^2 + 42^2) is your s

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

43.1740663*pi*10=?

OpenStudy (anonymous):

1 356.3533 so yes :)

OpenStudy (anonymous):

That's the same answer I got! Thank you so much!! :D

OpenStudy (anonymous):

de nada

OpenStudy (anonymous):

Isn't that the lateral surface area? Wouldn't the total surface area be 1671m2?

OpenStudy (anonymous):

@uri which of these are correct? They say its a lateral area but im confused

OpenStudy (anonymous):

Circular Cone Formulas in terms of radius r and height h: Volume of a cone: V = (1/3)πr2h Slant height of a cone: s = √(r2 + h2) Lateral surface area of a cone: L = πrs = πr√(r2 + h2) Base surface area of a cone (a circle): B = πr2 Total surface area of a cone: A = L + B = πrs + πr2 = πr(s + r) = πr(r + √(r2 + h2)) Circular Cone Calculations: Use the following additional formulas along with the formulas above. Given radius and height calculate the slant height, volume, lateral surface area and total surface area. Given r, h find s, V, L, A use the formulas above Given radius and slant height calculate the height, volume, lateral surface area and total surface area. Given r, s find h, V, L, A h = √(s2 - r2) Given radius and volume calculate the height, slant height, lateral surface area and total surface area. Given r, V find h, s, L, A h = (3 * v) / (πr2) Given radius and lateral surface area calculate the height, slant height, volume and total surface area. Given r, L find h, s, V, A s = L / (πr) h = √(s2 - r2) Given radius and total surface area calculate the height, slant height, volume and lateral surface area. Given r, A find h, s, V, L s = [A - (πr2)] / (πr) h = √(s2 - r2) Given height and slant height calculate the radius, volume, lateral surface area and total surface area. Given h, s find r, V, L, A r = √(s2 - h2) Given height and volume calculate the radius, slant height, lateral surface area and total surface area. Given h, V find r, s, L, A r = √[ (3 * v) / (π * h) ] Given slant height and lateral surface area calculate the radius, height, volume, and total surface area. Given s, L find r, h, V, A r = L / (π * s) h = √(s2 - r2)

OpenStudy (anonymous):

The formula for the surface area of a cone is SA=pi*r*l+B pi*10*43.174=1356.353295+the base which is 100*pi, the answer is D) about 1,671 I just took this test. :)